# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v1

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on Llama2 family models reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase Zhou et al. (2023a), LLMs Touvron et al. (2023); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Zhang et al. (2024c); Cohen et al. (2023); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Huang et al. (2023); Jiang et al. (2024b). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Therefore, it is essential to equip LLMs with new knowledge to them up-to-date.
@@
We introduce Wiki-Newpages-QA datasets to enable a comprehensive analysis of an LLM’s knowledge acquisition ability in terms of memorization, extraction, and reasoning.
We validate the efficacy of Self-Tuning on three crucial knowledge acquisition tasks using the Wiki-Newpages-QA datasets.
@@ 5.1 Setup
(2) Evaluation on Knowledge Retention We carry out assessments on two aspects of knowledge retention, specifically extraction and reasoning. For extraction, we evaluate the model’s performance in retaining factual knowledge using Natural Questions (NQ) Kwiatkowski et al. (2019) (i.e., , NQ-open Min et al. (2021)) and report EM and F1 scores. For reasoning, we assess the capability in retaining commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019) and report accuracy.
@@ 6 Conclusion
In this study, we introduce Self-Tuning, a framework aimed at enhancing an LLM’s ability to effectively learn from raw documents through self-teaching. Specifically, we create Self-Teaching, a self-supervised learning strategy that presents documents as plain texts along with various knowledge-intensive tasks derived directly from the documents, without the need for any mining patterns. Furthermore, we present three Wikipedia-Newpages-QA datasets to enable a comprehensive evaluation of an LLM’s knowledge acquisition capabilities, w.r.t., memorization, extraction, and reasoning across three distinct scenarios. Our findings show that Self-Tuning consistently yields superior performance on the knowledge acquisition tasks while demonstrating impressive knowledge retention performance. These results suggest the potential for broader applications of the proposed Self-Tuning, such as acquiring domain-specific knowledge.
@@
Ethics Statement
The primary objective of this study is to align with ethical standards that focus on mitigating hallucinations in LLMs. Throughout the research, we have consistently adhered to ethical guidelines.
During the Wiki-Newpages-QA data collection process using GPT-4, we carefully constructed prompts to eliminate any language that might discriminate against specific individuals or groups. These measures aimed to minimize potential negative effects on users’ well-being. Examples of these thoughtfully designed prompts can be found in Table 20 and Table 19. To further ensure dataset quality, we engaged student assistants to manually review the newly collected datasets. These datasets were confirmed to be of high quality, devoid of offensive content, false information, or any personally identifiable information. All student assistants involved were briefed on the research objectives before manual verification. Their privacy was guaranteed to be secure and confidential throughout the research process, and each participant received fair compensation.
The knowledge retention tasks utilized well-established benchmark datasets. Our study is dedicated to advancing knowledge while maintaining a strong commitment to privacy, fairness, and the well-being of all individuals and groups involved Zhou et al. (2023b).
@@ Appendix B Evaluation Results of Additional Compared Methods
Setup. To provide a comprehensive evaluation, we further verify the effectiveness of our proposed Self-Tuning by comparing it with three representative methods, i.e., standard instruction-tuning w/o forgetting, QA adaptation, and PIT, as shown in Table 8.
Results. We present the evaluation results in Table A. Our proposed Self-Tuning consistently outperforms the compared methods by a significant margin. Moreover, Self-Tuning substantially surpasses QA adaptation, suggesting the importance of equipping the model with the ability to systematically absorb knowledge from raw documents rather than merely providing it with question-answering skills.
@@ Appendix H Examination of QA Types in Open-ended Generation QA Datasets
We conduct an in-depth analysis of the QA types related to the factual information contained in the open-ended generation QA datasets, as shown in Table H.
@@ Appendix I Detailed Templates used in the Self-Teaching Strategy
The prompts used for constructing the QA datasets for open-ended generation and NLI tasks are presented in Table 18 and Table 19, respectively. The prompt used during the evaluation process is displayed in Table 20.

### v2

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on Llama2 family models reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††∗Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase Zhou et al. (2023), LLMs Touvron et al. (2023); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Zhang et al. (2024c); Cohen et al. (2023); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Huang et al. (2023); Jiang et al. (2024b). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Therefore, it is essential to equip LLMs with new knowledge to keep them up-to-date.
@@
We introduce Wiki-Newpages-2023-QA datasets to enable a comprehensive analysis of an LLM’s knowledge acquisition ability in terms of memorization, extraction, and reasoning.
We validate the efficacy of Self-Tuning on three crucial knowledge acquisition tasks using the Wiki-Newpages-2023-QA datasets.
@@ 5.1 Setup
(2) Evaluation on Knowledge Retention We carry out assessments on two aspects of knowledge retention, specifically extraction and reasoning. For extraction, we evaluate the model’s performance in retaining factual knowledge using Natural Questions (NQ) Kwiatkowski et al. (2019) (NQ-open Min et al. (2021)) and report EM and F1 scores. For reasoning, we assess the capability in retaining commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019) and report accuracy.
@@ 6 Conclusion
In this study, we introduce Self-Tuning, a framework aimed at enhancing an LLM’s ability to effectively learn from raw documents through self-teaching. Specifically, we create Self-Teaching, a self-supervised learning strategy that presents documents as plain texts along with various knowledge-intensive tasks derived directly from the documents, without the need for any mining patterns. Furthermore, we present three Wikipedia-Newpages-2023-QA datasets to enable a comprehensive evaluation of an LLM’s knowledge acquisition capabilities, w.r.t., memorization, extraction, and reasoning across three distinct scenarios. Our findings show that Self-Tuning consistently yields superior performance on the knowledge acquisition tasks while demonstrating impressive knowledge retention performance. These results suggest the potential for broader applications of the proposed Self-Tuning, such as acquiring domain-specific knowledge.
@@ Appendix B Evaluation Results of Additional Compared Methods
Setup. For a thorough assessment, we also examine the efficiency of our suggested Self-Tuning by contrasting it with four other notable methods, namely standard instruction-tuning without forgetting, QA adaptation, PIT, and mixed training, as displayed in Table 8.
Results. We present the evaluation results in Table A. Our Self-Tuning consistently shows superior performance, for example, it increases the EM by 11% on the knowledge extraction task.
In addition, Self-Tuning significantly outperforms QA adaptation in knowledge retention and extraction tasks, emphasizing the need to equip the model with the capability to systematically assimilate knowledge from raw documents instead of merely providing it with question-answering skills. Surprisingly, QA adaptation achieves high accuracy on the reasoning task. This could be attributed to an LLM’s ability to develop impressive reasoning skills during its pre-training phase, and a model with robust reasoning capabilities might excel in unfamiliar knowledge-intensive tasks Ovadia et al. (2024). Consequently, we regard knowledge extraction as the primary evaluation dimension, in line with previous studies Ovadia et al. (2024); Jiang et al. (2024b).
@@ Appendix H Examination of QA Types in Open-ended Generation QA Datasets
We perform a detailed analysis of the QA types associated with the factual information in the open-ended generation QA datasets, as displayed in Table H, by using the prompt in Table 21 with GPT-4.
@@ Appendix I Detailed Templates used in the Self-Teaching Strategy
The prompts used for constructing the QA datasets for open-ended generation and NLI tasks are presented in Table 18 and Table 19, respectively. The prompt used during the evaluation process is displayed in Table 20. The prompt used by GPT-4 for annotating QA types in the open-ended generation tasks of the Wiki-Newpages-2023-QA datasets is presented in Table 21.
@@
| The prompt used by GPT-4 for annotating QA types in the open-ended generation tasks of the Wiki-Newpages-2023-QA datasets |
| Below is a paragraph along with corresponding question and answer pairs. Your task is to analyze the paragraph and the question-answer pairs by categorizing the type of information they inquire about or provide. Use concise phrases to describe each category. For example: |
| Paragraph: <Andrew Turner (rugby union, born 2002) - Wikipedia> Andrew Turner (born 16 February 2002) is an English rugby union player, currently playing for the and . His preferred position is prop. |
| Question: When was Andrew Turner (rugby union, born 2002) born? |
| Answer: February 16, 2002. |
| Question: What nationality is Andrew Turner (rugby union, born 2002)? |
| Answer: English. |
| Question: What sport does Andrew Turner (rugby union, born 2002) play? |
| Answer: Rugby union. |
| Analysis: Types of question-answer pairs: (1) Birth date, (2) Nationality, (3) Sport/Profession. |
| Types of the paragraph: Biography - Biographical information about Andrew Turner, a rugby union player born in 2002, including his birth date, nationality, sport, and preferred position. |
| Below is a paragraph along with corresponding question and answer pairs. Your task is to analyze the paragraph and the question-answer pairs by categorizing the type of information they inquire about or provide. Use concise phrases to describe each category. For example: |
Paragraph: {paragraph}
{QA}
| Analysis: |